# A localized consensus-based sampling algorithm

### v2

@@ 5.2 Polarized consensus-based sampling
$\dot{U}_{t}=-(U_{t}-\tilde{\mu}_{\alpha,k}(\rho_{t};U_{t}))+\sqrt{2(\alpha+1)\,\tilde{\mathcal{C}}_{\alpha,k}(\rho_{t};U_{t})}\,\dot{W}_{t},$ (53)
with the polarized mean and covariance
@@
$\tilde{\mathcal{C}}_{\alpha,k}(\rho;u)$ $\coloneqq\frac{\int(v-\tilde{\mu}_{\alpha,k}(\rho;u))\otimes(v-\tilde{\mu}_{\alpha,k}(\rho;u))\,\tilde{w}_{\alpha,k}(v;u)\,\mathrm{d}\rho(v)}{\int\tilde{w}_{\alpha,k}(v;u)\,\mathrm{d}\rho(v)},$ (54b)
where
$\tilde{w}_{\alpha,k}(v;u)=k(v,u)\,\widehat{\pi}(v)^{\alpha}.$ (55)
@@
This choice of $k$ introduces a localization effect: the polarized mean $\tilde{\mu}_{\alpha,k}(\rho_{t};U_{t})$—the “target” of the drift in Eq. 53—tends to produce a value that is close to $U$, as well as having a large density in the posterior $\pi\propto\widehat{\pi}$. When $\pi$ is Gaussian, [4, Proposition 1] proves that $\pi$ is a stationary distribution of Eq. 53 with Eq. 56. More generally, a localized kernel attempts to lessen the Gaussian assumption present in CBS.
@@
    Localized CBS includes the correction term $\nabla_{u}\cdot\mathbf{C}_{U,t}$. As $\beta\rightarrow\infty$ and ${\kappa\rightarrow 0}$, then, we expect convergence to the corrected interacting Langevin dynamics, under which even non-Gaussian and multimodal $\pi$ are invariant. In polarized CBS, we do not expect a similar convergence.
@@
    By including the preconditioner $\mathbf{C}_{U,t}$ in the distance weighting for the localized mean, localized CBS has affine-invariant dynamics when $\mathbf{C}_{U,t}$ is a weighted or unweighted covariance. The polarization kernels proposed for polarized CBS in [4], on the other hand, are not affine-invariant.
@@
    Using $\widehat{\pi}$-based and distance-based weights in the (relatively cheap) mean computation is crucial in both methods, since that drives evolution towards high-probability regions and allows multimodal sampling. On the other hand, we find that the benefits of using weights in the (relatively expensive) covariance preconditioner are significantly less clear. Yet, this can be costly: distance-weighted statistics need to be computed for each of the $J$ particles, instead of once, globally. Polarized CBS uses the same weighting for both the mean and the covariance, while localized CBS decouples these two elements and allows to use a cheap, unweighted covariance as a preconditioner. Especially in higher-dimensional problems, this can make localized CBS the computationally cheaper method.
@@ 5.3 Localized ALDI
$\dot{U}_{t}=-\mathcal{C}^{G}(\rho_{t})\Gamma^{-1}\,(G(U_{t})-y)+\mathcal{C}(\rho_{t})\Gamma_{0}^{-1}\,(U_{t}-u_{0})+\sqrt{2\mathcal{C}(\rho_{t})}\,\dot{W}_{t},$ (57)
@@
Equation 57 approaches the sampling problem Eq. 1 explicitly in the context of a Bayesian inverse problem Eq. 2, estimating the parameters of a forward model $G$. It also assumes that $\pi_{\mathrm{prior}}=\mathcal{N}(u_{0},\Gamma_{0})$ and $\pi_{\mathrm{noise}}=\mathcal{N}(0,\Gamma)$ are Gaussians. In the case where $G$ is linear, Eq. 57 is equivalent to Eq. 8 and, hence, has $\pi$ as an invariant distribution [19]. For nonlinear $G$, this is generally not the case.
In [47], localization with a scalar $\lambda$ and positive definite matrix $D$ is added to Eq. 57, together with a correction term similar to the one in Eq. 14. This results in
$\dot{U}_{t}$ $=-\tilde{\mathcal{C}}_{\lambda D}^{G}(\rho_{t};U_{t})\Gamma^{-1}\,(G(U_{t})-y)-\tilde{\mathcal{C}}_{\lambda D}(\rho_{t};U_{t})\Gamma_{0}^{-1}\,(U_{t}-u_{0})$ (59)
$\qquad+\nabla_{u}\cdot\tilde{\mathcal{C}}_{\lambda D}(\rho_{t};U_{t})+\sqrt{2\tilde{\mathcal{C}}_{\lambda D}(\rho_{t};U_{t})}\,\dot{W}_{t},$
where
@@
$\tilde{\mathcal{C}}_{\lambda D}(\rho;u)$ $\coloneqq\frac{\int(v-\tilde{\mu}_{\lambda D}(\rho;u))\otimes(v-\tilde{\mu}_{\lambda D}(\rho;u))\,\tilde{w}_{\lambda D}(v;u)\,\mathrm{d}\rho(v)}{\int\tilde{w}_{\lambda D}(v;u)\,\mathrm{d}\rho(v)}$ (60b)
with
$\tilde{w}_{\lambda D}(v;u)=\exp\bigl{(}-(2\lambda)^{-1}\lVert v-u\rVert_{D}^{2}\bigr{)}.$ (61)
The definition of $\tilde{\mathcal{C}}_{\lambda D}^{G}$ is analogous. We stress again that this is a different concept of localization than the one discussed in Remark 1.1.
This localization improves performance for nonlinear $G$, making localized ALDI a viable method for our problem class. However, we point out two potential advantages of the localized CBS dynamics.
@@
    Similarly to polarized CBS, when the matrix $D$ in the distance-based weights of localized ALDI is fixed, the dynamics are not affine-invariant. In [47], affine-invariance is proven if $D=\mathcal{C}(\rho_{0})$, such that $D$ is also subject to any affine transformation. This can be considered a slightly weaker form of affine-invariance: in practice, the scaling of the parameters is often not fully known and $\rho_{0}$ cannot be chosen with optimal scaling. After all, if the scaling was known exactly, we could rescale the problem manually and would not need affine-invariant samplers. While the effect of the initial distribution may diminish over time in localized CBS (see, e.g., subsection 7.2 and Figure 3 in particular), choosing $D=\mathcal{C}(\rho_{0})$ in localized ALDI ensures that an incorrectly scaled initial distribution persists.
@@ 6 Discretizations and computational considerations
In this section, we bridge the gap between the theoretical localized CBS dynamics Eq. 29 and practical algorithms to solve problems of the form Eq. 1.
@@
$\widehat{\mathcal{C}}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i})\coloneqq\sum_{j=1}^{J}\omega_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}^{ij}\;(U^{j}-\widehat{\mu}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i}))\otimes(U^{j}-\widehat{\mu}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i}))$ (63)
@@
where $\omega_{\alpha,A}^{ij}=w_{\alpha,A}^{ij}/\sum_{k=1}^{J}w_{\alpha,A}^{ik}$ with $w_{\alpha,A}^{ij}=w_{\alpha,A}(U_{j};U_{i})$. In general, we denote the particle approximation to $\mathbf{C}(\rho;U^{i})$ by $\widehat{\mathbf{C}}(\bm{U};U^{i})$. As in the mean-field setting of section 4, we assume that this preconditioner is invertible. Notably, this requires that the $J$ particles span the entire space $\mathbb{R}^{d}$. While particle-based methods such as ALDI have been used with $J<d$ to sample in a linear subspace [20], we leave the application of localized CBS to those settings to future work.
@@
After particle discretization, we are left with a system of interacting SDEs. Since studying time discretizations is not a main point of focus of this paper, we will use the standard Euler–Maruyama discretization with $N$ timesteps of length ${\Delta\kern-0.8ptt}$ in our numerical tests, although others are possible [^6] Classical CBS uses a non-standard timestepping scheme, which helps them prove properties in the time-discrete setting. Polarized CBS adopts this same method. While EKS uses a linearly implicit split-step time discretization, both ALDI and localized ALDI use standard Euler–Maruyama.. We denote the $i$th particle at timestep $n$ by $U_{n}^{i}$ and the ensemble by $\bm{U}_{n}$.
@@
For the weighted mean in the drift term of localized CBS, we introduce a different particle approximation that forbids self-to-self interaction (which would otherwise have very large weights when particles are sparsely distributed) and randomly subsamples particles with a factor $\nu\in(0,1]$:
@@
Here, $\theta_{n}^{i,j}$ is uniformly random in $[0,1]$ and we write $\widehat{\mathbf{C}}\coloneqq\widehat{\mathbf{C}}(\bm{U}_{n};U_{n}^{i})$ for notational convenience. This idea is related to the random-batch method of [29], where interacting-particle systems are divided into small, random batches for a reduction in computational cost. When we use random batching, we will choose relatively large values of $\nu$ (e.g., $\nu=0.5$), so the cost difference is moderate; instead, we are motivated by our observation that random-batch interaction can improve sampling performance for higher-dimensional, multimodal problems. This is illustrated in section 7.
@@
By using $\widehat{\mathbf{C}}^{1/2}$ and simultaneously replacing the $d$-dimensional Brownian motion $W_{t}$ by an $n$-dimensional one, the Fokker–Planck equation remains the same. Henceforth, we define $\bar{U}\coloneqq\widehat{\mu}(\bm{U})$, $\bar{U}^{(i)}\coloneqq\widehat{\mu}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i})$, $w^{ij}\coloneqq w^{ij}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}$, and $\omega^{ij}\coloneqq\omega^{ij}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}$ for conciseness. The unweighted and weighted covariance have factors
@@
respectively. These are readily available and avoid computing the square root of $\widehat{\mathbf{C}}$. This should offer computational advantages unless $J\gg d$.
@@
In the weight function Eq. 24, localized CBS uses preconditioner-weighted squared norms ${\lVert U^{i}-U^{j}\rVert_{\widehat{\mathbf{C}}}^{2}=(U^{i}-U^{j})^{T}\widehat{\mathbf{C}}^{-1}(U^{i}-U^{j})}$. This quantity can be evaluated as follows: precompute
@@
The discretized dynamics contain the correction term $\nabla_{u^{i}}\cdot\widehat{\mathbf{C}}$, where $\widehat{\mathbf{C}}\coloneqq\widehat{\mathbf{C}}(\bm{U}_{n};U_{n}^{i})$. We now study this term for the preconditioners we consider in this article. When $\widehat{\mathbf{C}}=K$, clearly $\nabla_{u^{i}}\cdot\widehat{\mathbf{C}}=0$. When $\widehat{\mathbf{C}}=\widehat{\mathcal{C}}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U)$, the correction term is given in Lemma 6.1. The preconditioners $\widehat{\mathcal{C}}_{\alpha}(\bm{U})$ and $\widehat{\mathcal{C}}(\bm{U})$ correspond to Lemma 6.1 with $\lambda=\infty$ and, in the latter case, $\alpha=0$ as well. Note that the correction term contains a derivative of $\widehat{\pi}$ when $\alpha\neq 0$, which is problematic in our otherwise derivative-free method. In this paper we always choose $\alpha=0$. If situations are identified where $\alpha\neq 0$ is preferred, one could neglect the $\widehat{\pi}$ derivatives (see Remark 6.2) or construct a particle approximation (e.g., based on the proximal operator such as in section 3 or on the technique in [53]). This is left for future work.
@@
$\quad+\frac{1}{\lambda J}\bm{U}_{c}^{(i)}\,\bigl{(}\omega^{i}\odot\mathrm{diag}\bigl{(}(\bm{U}_{c}^{(i)})^{T}\widehat{\mathcal{C}}^{-1}(\bm{U}-U^{i}\bm{1}^{T})\bigr{)}\odot\bigl{(}(U_{c}^{i})^{T}\widehat{\mathcal{C}}^{-1}(\bm{U}-U^{i}\bm{1}^{T})\bigr{)}^{T}\bigr{)}$
@@
where $\odot$ denotes the Hadamard (elementwise) product, $\bm{U}_{c}\coloneqq\bm{U}-\bar{U}\bm{1}^{T}$, $\bm{U}_{c}^{(i)}\coloneqq\bm{(}\bm{U}-\bar{U}^{(i)}\bm{1}^{T})$, $U_{c}^{i}$ is the $i$th column of $\bm{U}_{c}$, $U_{c}^{(i),j}$ is the $j$th column of $\bm{U}_{c}^{(i)}$, $\widehat{\mathcal{C}}$ is a shorthand for $\widehat{\mathcal{C}}(\bm{U})$, and $\omega^{i}=(\omega^{i1},\ldots,\omega^{iJ})^{T}$.

### v3

@@ 5.2 Polarized consensus-based sampling
$\mathrm{d}U_{t}=-(U_{t}-\tilde{\mu}_{\alpha,k}(\rho_{t};U_{t}))\,\mathrm{d}t+\sqrt{2(\alpha+1)\,\tilde{\mathcal{C}}_{\alpha,k}(\rho_{t};U_{t})}\,\mathrm{d}W_{t},$ (51)
where we use the weights $\tilde{w}_{\alpha,k}(v;u)=k(v,u)\,\widehat{\pi}(v)^{\alpha}$ to define a polarized mean and covariance
@@
$\tilde{\mathcal{C}}_{\alpha,k}(\rho;u)$ $\coloneqq\frac{\int(v-\tilde{\mu}_{\alpha,k}(\rho;u))\otimes(v-\tilde{\mu}_{\alpha,k}(\rho;u))\,\tilde{w}_{\alpha,k}(v;u)\,\mathrm{d}\rho(v)}{\int\tilde{w}_{\alpha,k}(v;u)\,\mathrm{d}\rho(v)}.$ (52b)
@@
This choice of $k$ introduces a localization effect: the polarized mean $\tilde{\mu}_{\alpha,k}(\rho_{t};U_{t})$—the “target” of the drift in Eq. 51—tends to produce a value that is close to $U$, as well as having a large density in the posterior $\pi\propto\widehat{\pi}$. When $\pi$ is Gaussian and Eq. 53 is used for the kernel, [4, Proposition 1] proves that $\pi$ is a stationary distribution of Eq. 51 with Eq. 53. More generally, a localized kernel attempts to lessen the Gaussian assumption present in CBS in order to allow multimodal sampling.
@@
    Localized CBS includes the correction term $\nabla_{u}\cdot\mathbf{C}_{U,t}$. As $\beta\rightarrow\infty$ and $\kappa\rightarrow 0$}, we expect convergence to the interacting Langevin dynamics, under which even non-Gaussian and multimodal $\pi$ are invariant. In polarized CBS, we do not expect a similar convergence.
@@
    By including the preconditioner $\mathbf{C}_{U,t}$ in the distance weighting for the localized mean, localized CBS has affine-invariant dynamics when $\mathbf{C}_{U,t}$ is a weighted or unweighted covariance. The kernels proposed for polarized CBS in [4] are not affine-invariant as they weight the distance between two particles by a fixed matrix.
@@
    Using $\widehat{\pi}-$based and distance-based weights in the (relatively cheap) mean computation is crucial in both methods, since that drives evolution towards high-probability regions and allows multimodal sampling. On the other hand, we find that the benefits of using weights in the (relatively expensive) covariance preconditioner are significantly less clear. Yet, this can be costly: distance-weighted statistics need to be computed for each of the $J$ particles, instead of once, globally. Polarized CBS uses the same weighting for both the mean and the covariance, while localized CBS decouples these two elements and allows to use a cheap, unweighted covariance as a preconditioner. This can make localized CBS the computationally cheaper method in higher-dimensional problems.
@@ 5.3 Localized ALDI
$\mathrm{d}U_{t}=-\mathcal{C}^{G}(\rho_{t})\Gamma^{-1}\,(G(U_{t})-y)\,\mathrm{d}t+\mathcal{C}(\rho_{t})\Gamma_{0}^{-1}\,(U_{t}-u_{0})\,\mathrm{d}t+\sqrt{2\mathcal{C}(\rho_{t})}\,\mathrm{d}W_{t},$ (54)
@@
Equation 54 approaches the sampling problem Eq. 1 explicitly for a Bayesian inverse problem Eq. 2, estimating the parameters of a model $G$. It assumes that $\pi_{\mathrm{prior}}=\mathcal{N}(u_{0},\Gamma_{0})$} and $\pi_{\mathrm{noise}}=\mathcal{N}(0,\Gamma)$} are Gaussians. When $G$ is linear, Eq. 54 is equivalent to Eq. 8 and, hence, has $\pi$ as an invariant distribution [20]. For nonlinear $G$, this is generally not the case.
In [48], localization with a scalar $\lambda$ and positive definite matrix $D$ is added to Eq. 54, together with a correction term similar to the one in Eq. 14. This can be seen as Langevin diffusions where the gradients are approximated using the method in [60]. The resulting dynamics are
$\mathrm{d}U_{t}$ $=-\tilde{\mathcal{C}}_{\lambda D}^{G}(\rho_{t};U_{t})\Gamma^{-1}\,(G(U_{t})-y)\,\mathrm{d}t-\tilde{\mathcal{C}}_{\lambda D}(\rho_{t};U_{t})\Gamma_{0}^{-1}\,(U_{t}-u_{0})\,\mathrm{d}t$ (56)
$\qquad+\nabla_{u}\cdot\tilde{\mathcal{C}}_{\lambda D}(\rho_{t};U_{t})+\sqrt{2\tilde{\mathcal{C}}_{\lambda D}(\rho_{t};U_{t})}\,\mathrm{d}W_{t},$
where we use $\tilde{w}_{\lambda D}(v;u)=\exp(-(2\lambda)^{-1}\lVert v-u\rVert_{D}^{2})$ to define
@@
$\tilde{\mathcal{C}}_{\lambda D}(\rho;u)$ $\coloneqq\frac{\int(v-\tilde{\mu}_{\lambda D}(\rho;u))\otimes(v-\tilde{\mu}_{\lambda D}(\rho;u))\,\tilde{w}_{\lambda D}(v;u)\,\mathrm{d}\rho(v)}{\int\tilde{w}_{\lambda D}(v;u)\,\mathrm{d}\rho(v)}.$ (57b)
The definition of $\tilde{\mathcal{C}}_{\lambda D}^{G}$ is analogous. Localization in ALDI (which, we stress again, differs from the concept of localization discussed in Remark 1.1) improves performance for nonlinear $G$. However, we point out two potential comparative advantages of the localized CBS dynamics.
@@
    Similarly to polarized CBS, when the matrix $D$ in localized ALDI is fixed, the dynamics are not affine-invariant. In [48], affine-invariance is proven if $D=\mathcal{C}(\rho_{0})$, such that $D$ is also subject to any affine transformation. This can be considered a slightly weaker form of affine-invariance: in practice, the scaling of the parameters is often not fully known and $\rho_{0}$ cannot be chosen with optimal scaling. After all, if the scaling was known exactly, we could rescale the problem manually and would not need affine-invariant samplers. Whereas the effect of the initial distribution may diminish over time in localized CBS (see subsection 7.2 and Figure 3 in particular), choosing $D=\mathcal{C}(\rho_{0})$} in localized ALDI ensures that an incorrectly scaled initial distribution persists.
@@
5.4 Other methods
The localized CBS dynamics Eq. 28 arise from a specific approximation of the gradient in preconditioned Langevin dynamics, as detailed in section 3. Several other gradient-free Langevin-type variants have been proposed in the literature. Besides localized ALDI, discussed in subsection 5.3, notable examples include the following methods.
  1. (i)
    The method of [13] approximates gradients using only particles within a fixed neighborhood of the target particle. The authors of [13] report that this approach can become unstable because isolated particles may experience unrealistically large approximate gradients. To mitigate this, they propose evaluating the true gradient at such particles while the remaining ones use the ensemble-based approximation. Although conceptually interesting—and potentially adaptable to localized CBS—this algorithm is not fully gradient-free and is therefore applicable in a more limited setting than localized CBS.
  2. (ii)
    The multiscale sampler of [44] introduces two types of particles: a main (macro) particle that represent the distribution and a local swarm of auxiliary (micro) particles used to estimate gradients. This construction can handle non-Gaussian targets and may scale favorably with problem dimension because the particles stay tightly clustered. However, only the macro particle contributes to sampling; the computational effort invested in the auxiliary particles serves solely for gradient estimation. Moreover, since all particles in the swarm remain close, the method can explore only one mode at a time, which leads to metastability in multimodal problems. In contrast, every particle in localized CBS contributes to the sample ensemble, and the method can explore multiple modes simultaneously.
@@ 6 Discretizations and computational considerations
We now bridge the gap between the theoretical localized CBS dynamics Eq. 28 and practical algorithms to sample from Eq. 1.
@@
$\widehat{\mathcal{C}}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i})\coloneqq\sum\nolimits_{j=1}^{J}\omega_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}^{ij}\;(U^{j}-\widehat{\mu}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i}))\otimes(U^{j}-\widehat{\mu}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i}))$ (59)
@@
where $\omega_{\alpha,A}^{ij}=w_{\alpha,A}^{ij}/\sum_{k=1}^{J}w_{\alpha,A}^{ik}$ with $w_{\alpha,A}^{ij}=w_{\alpha,A}(U_{j};U_{i})$. In general, we denote the particle approximation to $\mathbf{C}(\rho;U^{i})$ by $\widehat{\mathbf{C}}(\bm{U};U^{i})$. As in the mean-field setting of section 4, we assume that this preconditioner is invertible. Notably, this requires that the $J$ particles span the entire space $\mathbb{R}^{d}$. Particle-based methods such as ALDI have been used with $J\ll d$ to sample in a linear subspace [21]; we leave the application of localized CBS to those settings to future work.
@@
Remark 6.1 (Scaling of the particle discretization to higher dimensions).
The performance of localized CBS in high dimensions has not been theoretically analyzed. Since the method relies on local interactions between particles distributed throughout parameter space for gradient estimations, increasing dimensionality may increase the typical distance between particles and hence negatively affect performance. Clarifying this requires dedicated study.
@@
After particle discretization, we are left with a system of interacting SDEs. Since time discretization is not a main point of focus of this paper, we use the standard Euler–Maruyama discretization with $N$ timesteps of length $\Delta\kern-0.8ptt$} in our numerical tests, although others are possible [^6] Classical CBS uses a non-standard timestepping scheme, which helps them prove properties in the time-discrete setting. Polarized CBS adopts this same method. While EKS uses a linearly implicit split-step time discretization, both ALDI and localized ALDI use standard Euler–Maruyama.. We denote the $i$th particle at timestep $n$ by $U_{n}^{i}$ and the ensemble by $\bm{U}_{n}$.
@@
For the weighted mean in the drift term of localized CBS, we introduce a particle approximation that forbids self-to-self interaction (which would have large weights when particles are sparse) and randomly subsamples particles with a factor $\nu\in(0,1]$:
@@
Here, $\theta_{n}^{i,j}$ is uniformly random in $[0,1]$ and we write $\widehat{\mathbf{C}}\coloneqq\widehat{\mathbf{C}}(\bm{U}_{n};U_{n}^{i})$ for notational convenience. This idea is related to the random-batch method of [30], where interacting-particle systems are divided into small, random batches for a reduction in computational cost. We will choose relatively large values of $\nu$ (e.g., $\nu=0.5$), so the cost difference is moderate; instead, we are motivated by our observation that random-batch interaction can improve sampling performance for higher-dimensional, multimodal problems. This is illustrated in section 7.
@@
By using $\widehat{\mathbf{C}}^{1/2}$ and simultaneously replacing the $d$-dimensional Brownian motion $W_{t}$ by an $n$-dimensional one, the Fokker–Planck equation remains the same. Henceforth, let us define $\bar{U}\coloneqq\widehat{\mu}(\bm{U})$$\bar{U}^{(i)}\coloneqq\widehat{\mu}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U^{i})$, $w^{ij}\coloneqq w^{ij}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}$, and $\omega^{ij}\coloneqq\omega^{ij}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}$ for conciseness. The unweighted and weighted covariance have factors
@@
respectively. These quantities are readily available and avoid computing the square root of $\widehat{\mathbf{C}}$. This should offer computational advantages unless $J\gg d$.
@@
Within the weight function Eq. 23, localized CBS uses squared norms $\lVert U^{i}-U^{j}\rVert_{\widehat{\mathbf{C}}}^{2}=(U^{i}-U^{j})^{T}\widehat{\mathbf{C}}^{-1}(U^{i}-U^{j})$} weighted by the preconditioner. This quantity can be evaluated as follows: precompute
@@
The discretized dynamics contain the correction term $^{\nabla_{u^{i}}\cdot\widehat{\mathbf{C}}}$, where we write $\widehat{\mathbf{C}}\coloneqq\widehat{\mathbf{C}}(\bm{U}_{n};U_{n}^{i})$}. We now study this term for the preconditioners we consider in this article. When $\widehat{\mathbf{C}}=K$, clearly $\nabla_{u^{i}}\cdot\widehat{\mathbf{C}}=0$; when $\widehat{\mathbf{C}}=\widehat{\mathcal{C}}_{\alpha,\lambda\widehat{\mathcal{C}}(\bm{U})}(\bm{U};U)$, the correction term is given in Lemma 6.2. The preconditioners $\widehat{\mathcal{C}}_{\alpha}(\bm{U})$ and $\widehat{\mathcal{C}}(\bm{U})$ correspond to Lemma 6.2 with $\lambda=\infty$} and, in the latter case, $\alpha=0$ as well. Note that the correction term contains a derivative of $\widehat{\pi}$ when $\alpha\neq 0$, which is problematic in our otherwise derivative-free method. In this paper we always choose $\alpha=0$. If situations are identified where $\alpha\neq 0$ is preferred, one could neglect the $\widehat{\pi}$ derivatives (see Remark 6.3) or construct a particle approximation (e.g., based on the proximal operator such as in section 3 or on the technique in [54]). This is left for future work.
@@
$\quad+\frac{1}{\lambda J}\bm{U}_{c}^{(i)}\,\bigl(\omega^{i}\odot\mathrm{diag}\bigl((\bm{U}_{c}^{(i)})^{T}\widehat{\mathcal{C}}^{-1}(\bm{U}-U^{i}\bm{1}^{T})\bigr)\odot\bigl((U_{c}^{i})^{T}\widehat{\mathcal{C}}^{-1}(\bm{U}-U^{i}\bm{1}^{T})\bigr)^{T}\bigr)\hskip-8.5359pt$
@@
where we denote by $\odot$ the Hadamard (elementwise) product. Furthermore, $\bm{U}_{c}\coloneqq\bm{U}-\bar{U}\bm{1}^{T}$ and $\bm{U}_{c}^{(i)}\coloneqq\bm{(}\bm{U}-\bar{U}^{(i)}\bm{1}^{T})$}. We write $U_{c}^{i}$ for the $i$th column of $\bm{U}_{c}$, $U_{c}^{(i),j}$ for the $j$th column of $\bm{U}_{c}^{(i)}$, $\widehat{\mathcal{C}}$ as a shorthand for $\widehat{\mathcal{C}}(\bm{U})$, and $\omega^{i}=(\omega^{i1},\ldots,\omega^{iJ})^{T}$.